A pedestrian is running at his maximum speed of 6.0 m/sec to catch a bus stopped at a traffic light. When his is 15m from the bus, the light changes and the bus accelerates uniformly at 1m/s/s. Does he make it to the bus? If so , how far does he have to run in order to catch it? If not, how close does he get?
1
Expert's answer
2019-06-10T09:41:14-0400

The position of pedestrian is described by:


"x_p=v_0t=6t"

The position of the bus:


"x_b=15+\frac{at^2}{2}=15+0.5t^2"

At the moment when they meet:


"x_p=x_b"

"6t=0.5t^2+15"

The smallest positive solution of this equation is:

"t=3.55s"

So, at t=3.55s, the pedestrian catches the bus.

The distance is:

"S=v_0t=6*3.55= 21.3 m"